Find the x-intercepts of the parabola y=x squared + 2x -15.
Answer by stanbon(48535)

Find the x-intercepts of the parabola y=x squared + 2x -15.
-----------
Let y=0 to find the x-intercepts, as follows:
x^2+2x-15=0
(x+5)(x-3)=0
x=3 or x=-5
These are the x-intercepts.
cheers,
Stan H.